One side of a rectangular patio is 6 feet longer than the other. If the area of the patio is 91 square feet, what are the dimensions of the patio?
Answer by edjones(8007)

x, x+6 the two sides of the patio.
x(x+6)=91
x^2+6x=91
x^2+6x-91=0
(x+13)(x-7)=0
x=7', x+6=13' The two sides.
.
Ed